solve each radical equation by using algebra.
square root of 2x+1 = square root of 4x-4

Answer by stanbon(75887)

sqrt(2x+1)=sqrt(4x-4)
Square both sides to get:
2x+1=4x-4
2x=5
x=5/2
Cheers,
Stan H.